# Beyond Ensembles: Simulating All-Atom Protein Dynamics in a Learned Latent Space

### v3

@@ 4.2 Long-horizon stability: autoregressive NN is most robust
Figure 2 tracks RMSD/lDDT [^2] The local Distance Difference Test (lDDT) evaluates local structural quality by measuring the preservation of inter-atomic distances. over rollout length and reports the failure time (first frame with lDDT <0.65<0.65 relative to the initial frame). On alanine dipeptide, Koopman and the autoregressive NN remain stable for thousands of frames (failure at 4,443 and 3,176, respectively), whereas Langevin terminates early (206). On $_{\mbox{1}}A$R, the NN completes the full 10,000-frame horizon without failure; Langevin reaches 7,476 and Koopman 5,740. Taken together, the NN is the most robust propagator for long-horizon rollouts on both a small peptide and a large GPCR, with Koopman providing a competitive linear baseline and Langevin exhibiting sensitivity to score/integration settings.
@@
We attribute the early failure on alanine to three factors: (i) a sparsely sampled latent manifold from a large MD frame stride, yielding a rough score estimate $s(z)=\nabla_{z}\log p(z)$; (ii) larger apparent per-frame displacements in a tiny system (RMSD $\sim$0.9 Å) making the lDDT threshold stringent; and (iii) step sizes tuned for GPCRs that are aggressive for alanine under an imperfect score, pushing the SDE off-manifold. In $_{\mbox{1}}A$R, motions distribute over many atoms (lDDT penalizes less), and subsampling (s=20s{=}20) further smooths per-frame jumps.
@@ 4.3 Fidelity to the Equilibrium Ensemble
Table 2: Dihedral free-energy fidelity (JSD; lower is better). Values computed on normalized 2D densities with periodic binning.
@@
On backbone dihedrals, the autoregressive NN consistently outperforms the other propagators. For alanine dipeptide, the NN achieves the lowest divergence (JSD =0.061=0.061), accurately capturing the canonical Ramachandran basins; the Koopman model’s linear rigidity yields a markedly less accurate landscape (JSD =0.138=0.138). This trend is even more pronounced on $_{\mbox{1}}A$R, where the NN again produces the most faithful backbone surface (JSD =0.019=0.019), while both Koopman and Langevin lag behind.
@@
The picture reverses for side-chain rotamers on $_{\mbox{1}}A$R. Here, the score-guided Langevin propagator is superior, yielding the sharpest rotamer bands and the lowest JSD ($0.058$). This likely reflects its thermodynamically consistent drift (Ts(z)T\,s(z)) plus isotropic noise, which excels at sampling sharp, multi-modal distributions corresponding to distinct rotameric states. The NN remains close (JSD =0.067=0.067) but slightly oversmooths high-barrier regions, whereas Koopman merges them.
Taken together, these results highlight a trade-off: the autoregressive NN excels at capturing broad backbone dynamics, while the score-guided Langevin propagator is superior for resolving fine-grained, multi-modal side-chain states.
@@ 4.4 GPCR function: activation surface in $_{\mbox{2}}A$R
Figure 4: Functional free–energy surface for $_{\mbox{2}}A$R. Left: 2D free energy over TM3–6 and TM3–7 distances (reference heatmap; model contours overlaid). Right: 1D profiles along each coordinate. Langevin spans the diagonal valley and its shoulders; the autoregressive NN follows the same valley with a tighter footprint; Koopman centers the basin but amplifies apparent barriers and reduces anisotropy. Free-energy surfaces for each propagator are averaged over 5 independent replicates.
@@
Along TM3–7 (top-right), Langevin and the NN track the position of the principal minimum and its curvature, with Langevin retaining broader support across the valley shoulders. Koopman exhibits a narrow well and steeper rise away from the minimum, consistent with its tendency to compress variance on the 2D map. Along TM3–6 (bottom-right), Langevin again follows the reference profile across both flanks, while the NN stays close near the minimum but with a slightly skewed tail. Koopman produces a compact basin and elevated shoulders, which reduces sampling of the transition corridor toward larger TM3–6 distances limiting its exploration of the activation corridor and showing that the non-linear propagators are more effective at tracing the complete activation pathway.
4.5 Outlook
Our comparison reveals a fundamental trade-off between the stability of data-driven models and the fine-grained physical accuracy of score-guided methods. This points toward several promising future directions. First is the development of hybrid propagators that could leverage neural networks for stable, long-horizon integration of slow collective motions while using a score-guided Langevin update for local, high-frequency side-chain sampling to ensure thermodynamic correctness. Second, a key avenue for improvement is the end-to-end training of the encoder and propagator to discover a latent representation optimized for stable temporal evolution. Such a latent space could be made more robust by enforcing thermodynamic consistency constraints or by explicitly training the model with a denoising objective, where it learns to predict the original protein structure from an input dataset to which noise has been added. For greater realism, models require rigorous kinetic calibration to match simulation or experimental timescales and should incorporate stronger physical priors like symmetry-aware (SE(3)) architectures, though their computational cost remains a significant challenge for large proteins.
@@ A.2 LD-FPG based Encoder-Decoder Backbone
Table 3: Configuration of the Pre-trained LD-FPG Backbone Model.
@@
Encoder (HNO)
@@
   Pooling Output Size ($H\times W$) $50\times 2$
   Flattened Latent Dimension ($d_{\text{flat}}$) 100  (from $50\times 2$)
@@ A.6 Contact map Pearson correlation
A.7 Ablation Studies on $_{\mbox{1}}A$R
@@
A.7.1 (i) Encoder per-atom width $d_{z}\in\{8,16,32\}$
@@
A.7.2 (ii) Pooling width $H$ with W=2W{=}2 (so $d_{\text{flat}}{=}2H$)
@@
A.7.3 (iii) Denoiser/Langevin settings
@@
A.8 Supplementary Figures
@@
A.9 Supplementary Tables

### v4

@@ 4.2 Long-horizon stability: autoregressive NN is most robust
Figure 2 tracks RMSD/lDDT [^2] The local Distance Difference Test (lDDT) evaluates local structural quality by measuring the preservation of inter-atomic distances. over rollout length and reports the failure time (first frame with lDDT <0.65<0.65 relative to the initial frame). This threshold was empirically calibrated: in our analysis, trajectories dropping below 0.65 consistently exhibited unphysical degradation (e.g., secondary structure melting or steric clashes), distinguishing true model failure from valid conformational drift (where lDDT typically remains >0.70>0.70). On alanine dipeptide, the autoregressive NN remains stable for the full 10,000-frame horizon without failure. The Koopman operator persists for thousands of frames (failure at 4,443), whereas Langevin terminates early (206) due to the high sensitivity of the score estimate on this small system. On $_{\mbox{1}}A$R, the NN again completes the full 10,000-frame horizon without failure; Langevin reaches 7,476 and Koopman 5,740. Taken together, the NN is the most robust propagator for long-horizon rollouts on both a small peptide and a large GPCR, with Koopman providing a competitive linear baseline and Langevin exhibiting sensitivity to score/integration settings.
@@
We attribute the early failure on alanine to three factors: (i) a sparsely sampled latent manifold from a large MD frame stride, yielding a rough score estimate $s(z)=\nabla_{z}\log p(z)$; (ii) larger apparent per-frame displacements in a tiny system (RMSD $\sim$0.9 Å) making the lDDT threshold stringent; and (iii) step sizes were calibrated to match the reference MD’s RMSD-per-frame. For the fast-moving alanine system, this protocol necessitated a large integration step ($\Delta t=10^{-8}$), which proved numerically aggressive for the Langevin integrator compared to the conservative step ($\Delta t=10^{-10}$) derived for the slower $_{\mbox{1}}A$R system, pushing the SDE off-manifold.
@@ 4.3 Fidelity to the Equilibrium Ensemble
Table 2: Dihedral free-energy fidelity (JSD; lower is better). Values represent the mean $\pm$ standard deviation over 5 independent runs. Bold indicates the best mean performance.
Propagator ADP $(\phi,\psi)$ $_{\mbox{1}}A$R $(\phi,\psi)$ $_{\mbox{1}}A$R $(\chi_{1},\chi_{2})$
Koopman $138\pm 0.015 0.0$$043\pm 0.006$ $0.121\pm 0.012$
Autoreg. NN $\mathbf{0.061\pm 0.003}$ $\mathbf{0.019\pm 0.001}$ $0.067\pm 0.002$
@@
On backbone dihedrals, the Autoregressive NN consistently outperforms the other propagators in both accuracy and stability. For alanine dipeptide, the NN achieves the lowest divergence (JSD $0.061\pm 0.003$), accurately capturing the canonical Ramachandran basins with high reproducibility. In contrast, the Koopman model’s rigidity yields a less accurate landscape (JSD =0.138=0.138) with noticeable variance across runs ($\pm 0.015$). This trend is even more pronounced on $_{\mbox{1}}A$R, where the NN produces the most faithful backbone surface (JSD =0.019=0.019) and negligible variance ($\pm 0.001$), confirming it as the most robust estimator for global backbone topology.
@@
The picture reverses for side-chain rotamers on $_{\mbox{1}}A$R. Here, the score-guided Langevin propagator achieves the best mean performance (JSD =0.058=0.058), yielding the sharpest rotamer bands. This likely reflects its thermodynamically consistent drift (Ts(z)T\,s(z)) plus isotropic noise, which excels at sampling sharp, multi-modal distributions. However, this comes at the cost of higher variance ($\pm 0.009$) compared to the NN ($\pm 0.002$), as the stochastic nature of the integrator leads to greater fluctuation between independent trajectories. The NN remains competitive (JSD =0.067=0.067) and highly stable, but slightly oversmooths high-barrier regions compared to the Langevin baseline.
Taken together, these results highlight a three-way trade-off: the Autoregressive NN is the most stable and accurate for broad backbone dynamics; the Score-Guided Langevin propagator is superior for resolving fine-grained, multi-modal side-chain states albeit with higher stochastic variance; and Koopman provides a lightweight but more rigid baseline with moderate variability.
@@ 4.4 GPCR function: activation surface in $_{\mbox{2}}A$R
Figure 4: Functional free–energy surface for $_{\mbox{2}}A$R. (Top Left) Reference MD free energy heatmap projected onto TM3–6 and TM3–7 distances, showing the transition valley. (Center & Bottom-Left) Generated ensembles for the Autoregressive NN (cyan), Langevin (orange), and Koopman (purple) propagators, plotted as contours over the reference background (grey). (Right Column) 1D free energy profiles projected along the TM3–7 (top) and TM3–6 (bottom) reaction coordinates.
@@
The 1D free energy profiles (Fig. 4, right column) quantify these deviations. Along TM3–7 (top-right), Langevin and the NN track the position of the principal minimum and its curvature, with Langevin retaining broader support across the valley shoulders. Koopman exhibits a narrow well and steeper rise away from the minimum, consistent with its tendency to compress variance on the 2D map. Along TM3–6 (bottom-right), Langevin again follows the reference profile across both flanks, while the NN stays close near the minimum but with a slightly skewed tail. Koopman produces a compact basin and elevated shoulders, which reduces sampling of the transition corridor toward larger TM3–6 distances limiting its exploration of the activation corridor and showing that the non-linear propagators are more effective at tracing the complete activation pathway.
5 Outlook
Our comparison reveals a fundamental trade-off between the stability of data-driven models and the fine-grained physical accuracy of score-guided methods. This points toward several promising future directions. First, motivated by the complementary strengths observed in Table 2 (autoregressive NN for backbone, Langevin for side-chains), we envision hybrid propagators that leverage neural networks for the stable, long-horizon integration of slow collective motions, while employing a score-guided Langevin update for local, high-frequency sampling to ensure thermodynamic correctness. Second, while our TICA analysis confirms that GLDP captures physical timescales and memory effects on complex systems where baselines fail, future work could enforce explicit kinetic constraints during training, such as spectral matching losses, to further refine this fidelity. Third, addressing the need for transferability, moving beyond a fixed backbone to the end-to-end training of the encoder and propagator on multi-protein datasets could discover latent representations optimized for generalizable temporal evolution, potentially bridging the gap between generalist foundation models and the high thermodynamic fidelity required for specific drug targets.
Reproducibility Statement.
We release a repository with all source code for the encoder–propagator–decoder pipeline, training, and evaluation, plus fixed random seeds and an environment file to pin dependencies. The main paper outlines the pipeline (Section 3) and the evaluation metrics (Section 4.3). Full implementation details appear in the Appendix: datasets and preprocessing (Appendix A.1), backbone configuration (Appendix A.2), propagators and hyperparameters (Appendix A.3, A.4, A.5), and ablations (Appendix A.8). The repository hosts the preprocessed latent trajectories used in our runs and links to the raw MD sources. To aid inspection, representative structure snapshots for ADP, 7JFL, $_{\mbox{1}}A$R, and $_{\mbox{2}}A$R are included in the Appendix (Figure 5), while short videos of the rollouts are available at the anonymous repository link.
@@ A.1 Simulation Datasets and Code availability
HIV-1 Protease (1R6W).
To evaluate performance on mixed secondary structures ($\alpha$-helices and $\beta$-sheets), we utilized the HIV-1 Protease system (PDB: 1R6W). We sourced the MD trajectory from the ATLAS databse dsimb.inserm.fr/ATLAS/…/1r6W.
@@ A.2 LD-FPG based Encoder-Decoder Backbone
Table 3: Configuration of the Pre-trained LD-FPG Backbone Model. Note: Pooling is performed immediately after the graph convolutions to form the latent space.
@@
Encoder & Pooling (Latent Construction)
   Graph Type Chebyshev GNN (HNO)
@@
   Pooling Output Size (Latent Bottleneck) 100
Decoder (Reconstruction)
@@ A.6 Contact map Pearson correlation
A.7 Methodology for Kinetic Comparison
To quantitatively evaluate the kinetic properties of the generated molecular dynamics trajectories against the ground truth simulations, we employed a two-stage analysis based on Time-lagged Independent Component Analysis (TICA) and Autocorrelation Functions (ACF). This approach allows us to identify the slowest collective motions within the system and measure their characteristic timescales.
A.7.1 Step 1: Featurization
The high-dimensional trajectory data was first transformed into a feature space suitable for kinetic analysis. We used the pairwise distances between all C-alpha atoms as our features. This choice of internal coordinates ensures that the analysis is invariant to rotational and translational motions of the protein as a whole. For a protein with $N$ C-alpha atoms, this results in a feature vector $x(t)\in\mathbb{R}^{N(N-1)/2}$ for each time frame $t$.
A.7.2 Step 2: Time-lagged Independent Component Analysis (TICA)
TICA is a powerful dimensionality reduction technique that finds the slowest, most dynamically relevant degrees of freedom from a molecular dynamics simulation. The objective of TICA is to find a linear combination of input features that maximizes the autocorrelation of the resulting components at a specified lag time, $\tau$. This is achieved by solving the generalized eigenvalue problem:
$C(\tau)v_{i}=\lambda_{i}C(0)v_{i}$ (6)
where:
  • $C(0)$ is the standard covariance matrix of the feature vector $x(t)$:
    $C(0)=\mathbb{E}\left[(x(t)-\mu)(x(t)-\mu)^{T}\right]$ (7)
  • $C(\tau)$ is the time-lagged covariance matrix at a lag time $\tau$:
    $C(\tau)=\mathbb{E}\left[(x(t)-\mu)(x(t+\tau)-\mu)^{T}\right]$ (8)
  • $\mu=\mathbb{E}[x(t)]$ is the mean of the feature vector.
  • $\lambda_{i}$ are the eigenvalues, which represent the autocorrelation of the $i$-th component at lag time $\tau$.
  • $v_{i}$ are the eigenvectors, which are the TICA components (TICs) or collective coordinates.
The TICA model was fit exclusively on the ground truth trajectory data to define the principal slow kinetic modes of the system.
A.7.3 Step 3: Calculation of Implied Timescales
The eigenvalues $\lambda_{i}$ obtained from the TICA decomposition are directly related to the characteristic "implied timescale" $t_{i}$ of each component. This timescale represents the decorrelation time of the corresponding TIC. The relationship is given by:
$t_{i}=-\frac{\tau}{\ln|\lambda_{i}|}$ (9)
The largest eigenvalue, $\lambda_{1}$, corresponds to the slowest process in the system, and its associated implied timescale, $t_{1}$, provides a quantitative measure of the duration of this dominant kinetic event. We report this slowest timescale as a key metric for kinetic validity.
A.7.4 Step 4: Projection onto the Slowest Mode
Once the TICA model was trained on the ground truth data, both the ground truth and the generated trajectories were projected onto the first (slowest) independent component, TIC1 (corresponding to eigenvector $v_{1}$). This projection reduces the high-dimensional conformational dynamics to a one-dimensional time series, $z(t)$, that captures the system’s evolution along its most significant slow coordinate.
$z(t)=v_{1}^{T}(x(t)-\mu)$ (10)
A.7.5 Step 5: Autocorrelation and Decorrelation Time
To provide a more direct comparison of the system’s "memory" along the slowest mode, we computed the normalized autocorrelation function (ACF) of the time series $z(t)$ for both trajectories. The ACF at a time lag $\Delta t$ is defined as:
$A(\Delta t)=\frac{\mathbb{E}\left[(z(t)-\bar{z})(z(t+\Delta t)-\bar{z})\right]}{\sigma_{z}^{2}}$ (11)
where $\bar{z}$ is the mean and $\sigma_{z}^{2}$ is the variance of the time series $z(t)$. The decorrelation time, $t_{decorr}$, is formally defined as the time lag at which the ACF decays to $1/e$.
$A(t_{decorr})=\frac{1}{e}\approx 0.368$ (12)
This value quantifies how long it takes for the system’s conformation along TIC1 to become statistically independent of its initial state. A close match in the decorrelation times between the ground truth and the generated trajectory indicates high kinetic fidelity.
A.8 Ablation Studies on $_{\mbox{1}}A$R
@@
A.8.1 (i) Encoder per-atom width $d_{z}\in\{8,16,32\}$
@@
A.8.2 (ii) Pooling width $H$ with W=2W{=}2 (so $d_{\text{flat}}{=}2H$)
@@
A.8.3 (iii) Denoiser/Langevin settings
@@
A.9 Ablation Studies on Data Efficiency and Temporal Resolution
To assess the robustness of the Graph Latent Dynamics Propagator (GLDP) beyond the standard benchmark settings, we performed targeted ablations on the 7JFL system. Specifically, we investigate the dependence of the Autoregressive NN propagator on training data volume, the impact of increasing the temporal prediction stride (frame skipping), and the benefit of residual parameterization.
A.9.1 Dependence on Trajectory Length (Data Efficiency)
We trained the Autoregressive NN propagator on subsets of the full 7JFL MD trajectory, ranging from 5% to 100% of the available frames. We evaluated structural fidelity via Backbone Dihedral JSD and kinetic fidelity via the slowest TICA timescale ($t_{1}$). We also report the Validation MSE, defined as the mean squared error between the predicted latent state $\hat{z}_{t+1}$ and the ground truth $z_{t+1}$ on a held-out test set, to quantify the local predictive accuracy of the propagator.
As shown in Table 10, the model exhibits a sharp performance threshold. At extremely low data regimes (5%), the propagator fails to learn the underlying physics, resulting in a collapse of the slow dynamics ($t_{1}\approx 22$) and poor structural recovery. While structural equilibrium (JSD) begins to stabilize around 25% data usage.
Table 10: Data Efficiency on 7JFL (Autoregressive NN). Structural and kinetic metrics as a function of training data fraction. (Mean of 3 runs).
Data Fraction Validation MSE ($10^{-4}$) Backbone JSD ($\downarrow$) TICA $t_{1}$ (steps)
5% 25.5 0.145 22.1
10% 7.2 0.065 95.4
25% 4.8 0.043 188.2
50% 4.1 0.039 215.6
75% 3.8 0.038 223.8
100% (Main Paper) 3.7 0.037 226.4
A.9.2 Impact of Frame Skipping (Temporal Stride)
Standard training minimizes the one-step error $z_{t}\to z_{t+1}$. To test the model’s capability to model coarser timescales, we trained separate models to predict $z_{t}\to z_{t+k}$ for stride $k\in\{1,\dots,10\}$.
Table 11 reports the validation MSE and the TICA timescale ($t_{1}$) in simulation steps. To compare kinetics fairly, we calculate the "Implied Physical Time" ($t_{1}\times k$). We observe a systematic reduction in physical memory as the stride increases. For $k=1$ to $k=3$, the physical timescale remains robust ($\approx 200+$ units). However, at larger strides ($k=10$), the implied physical time drops significantly. This indicates that predicting over large temporal gaps may cause the network to "smooth out" energy barriers, leading to accelerated transition kinetics and a loss of long-term memory.
Table 11: Effect of Frame Skipping on 7JFL. Validation MSE and Kinetic Fidelity. Note that $t_{1}$ is reported in simulation steps; Implied Physical Time is $t_{1}\times k$.
Stride ($k$) Validation MSE ($10^{-4}$) TICA $t_{1}$ (steps) Implied Phys. Time ($t_{1}\times k$)
1 (Baseline) 3.7 226.4 226.4
2 3.9 108.5 217.0
3 4.2 68.2 204.6
4 4.8 46.1 184.4
5 6.5 30.5 152.5
10 15.2 8.4 84.0
A.9.3 Residual vs. Direct Prediction
Finally, we addressed the architectural choice of predicting the next state directly, $z_{t+1}=f_{\theta}(z_{t})$ (as used in the main paper), versus predicting the residual, $z_{t+1}=z_{t}+f_{\theta}(z_{t})$.
As detailed in Table 12, both architectures converge to nearly identical performance. The residual formulation achieves a marginally lower validation MSE ($3.6$ vs $3.7$), but the resulting physical kinetics (TICA $t_{1}$) are statistically indistinguishable ($228$ vs $226$). This confirms that while residual learning is a valid optimization, the Direct Prediction formulation used in our main results is sufficient to capture the transition dynamics accurately.
Table 12: Architectural Comparison on 7JFL. Validation error and kinetic fidelity.
Architecture Validation MSE ($10^{-4}$) TICA $t_{1}$ (steps)
Direct Prediction (Main Paper) 3.7 226.4
Residual Prediction 3.6 228.1
A.9.4 Computational Efficiency of Latent Propagators
To quantify the utility of each propagator for long-horizon simulations, we benchmarked the wall-clock time required to generate trajectories of varying lengths. This analysis was performed for the 7JFL protein system on a single NVIDIA L40S GPU. The results clarify the trade-off between physical accuracy and raw speed.
Table 13: Wall-Clock Time for Trajectory Generation on 7JFL (L40S GPU)
Number of Steps Koopman (Average) Autoregressive NN Score-Guided Langevin
1,000 0.15s 2.62s 2.17s
10,000 0.19s 3.42s 4.03s
100,000 0.59s 12.53s 23.88s
A.10 Supplementary Figures
@@
Figure 6: Alanine Dipeptide Backbone Free Energy. Comparison of the ($\phi,\psi$) free energy surface for (i) Ground Truth MD, (ii) LSS, (iii) GeoTDM, and (iv) GLDP (Autoregressive NN).
Figure 7: $_{\mbox{1}}A$R Backbone Free Energy. Comparison of aggregated backbone ($\phi,\psi$) distributions. LSS (ii) shows a noisy, high-entropy landscape. GLDP (iv) and MD-Gen (iii) successfully recover the tight $\alpha$-helical basins observed in the Ground Truth (i).
Figure 8: $_{\mbox{1}}A$R Side-chain Free Energy. Comparison of aggregated side-chain ($\chi_{1},\chi_{2}$) distributions. The ability to capture rotameric packing is critical for stability. LSS (ii) fails to resolve clear rotamer wells. GLDP (iv) accurately reproduces the multi-modal rotameric density seen in the Ground Truth (i).
A.11 Supplementary Tables